# M4GT-Bench: Evaluation Benchmark for Black-Box Machine-Generated Text Detection Аннотация

### v1

@@
M4GT-Bench: Evaluation Benchmark for Black-Box
Machine-Generated Text Detection
@@
M4GT-Bench: Evaluation Benchmark for Black-Box
Machine-Generated Text Detection
Yuxia Wang,† Jonibek Mansurov,†, Petar Ivanov,† Jinyan Su,† Artem Shelmanov,† Akim Tsvigun,† Osama Mohammed Afzal,† Tarek Mahmoud,† Giovanni Puccetti,§ Thomas Arnold,¶ Alham Fikri Aji,† Nizar Habash,†‡ Iryna Gurevych,† Preslav Nakov† †Mohamed bin Zayed University of Artificial Intelligence, UAE ¶TU Darmstadt, Germany §Institute of Information Science and Technology, Italy ‡New York University Abu Dhabi, UAE {yuxia.wang, jonibek.mansurov, preslav.nakov}@mbzuai.ac.ae
@@
M4GT-Bench: Evaluation Benchmark for Black-Box
Machine-Generated Text Detection
Yuxia Wang,† Jonibek Mansurov,†, Petar Ivanov,† Jinyan Su,† Artem Shelmanov,† Akim Tsvigun,† Osama Mohammed Afzal,† Tarek Mahmoud,† Giovanni Puccetti,§ Thomas Arnold,¶ Alham Fikri Aji,† Nizar Habash,†‡ Iryna Gurevych,† Preslav Nakov† †Mohamed bin Zayed University of Artificial Intelligence, UAE ¶TU Darmstadt, Germany §Institute of Information Science and Technology, Italy ‡New York University Abu Dhabi, UAE {yuxia.wang, jonibek.mansurov, preslav.nakov}@mbzuai.ac.ae
@@ 1 Introduction
Task 2: Multi-Way Machine-Generated Text’s Generator Detection. This task challenges the identification of the specific language model responsible for generating a given full text. The text may be either human-written or crafted by a particular LLM. A different perspective to look at task 2 is that we are attributing authorship to a specific model Munir et al. (2021).
Task 3: Human-Written to Machine-Generated Text Change Point Detection. In this task, the goal is to precisely estimate the boundary within the mixed text, where the transition occurs from human-written to machine-generated content.
@@
We introduce a novel task formulation for MGT detection, where human-machine mixed text is explored. It is the first attempt to evaluate the ability to detect a boundary between human-written and machine-generated texts.
We construct the most diverse public multilingual MGT detection benchmark involving nine languages, six domains, nine LLM generators and three different tasks.
Human evaluation utilizing few-shot learning, yields a performance level below random chance in distinguishing different LLM generations, suggesting the challenges of task 2. We further demonstrate strong baselines for each task and investigate their performance in various experimental settings.
@@ 2 Related Work
The task of detecting machine-generated text has long been formulated as a binary classification problem Zellers et al. (2019); Gehrmann et al. (2019a); Solaiman et al. (2019); Ippolito et al. (2019), where the goal is to distinguish texts generated by certain language models from those authored by humans. However, with the rapid advancement of a wide range of LLMs such as GPT-4, Bard, Cohere and Claude, they can be easily used to generate texts through simple API calls. There is a growing need for more fine-grained classification that can not only identify the nature of the texts (i.e., whether it is machine-generated or human written), but also its specific source (i.e., which LLM generates it?). This is also known as authorship attribution Uchendu et al. (2020); Venkatraman et al. (2023); Rivera Soto et al. (2024).
@@
In general, there are mainly two types of machine-generated text detector: supervised approaches Wang et al. (2023); Uchendu et al. (2021); Zellers et al. (2019); Zhong et al. (2020); Liu et al. (2022) and unsupervised methods based on white-box features such as likelihood and log-rank Solaiman et al. (2019); Ippolito et al. (2019); Mitchell et al. (2023); Su et al. (2023); He et al. (2023); Hans et al. (2024), and watermarking Kirchenbauer et al. (2023); Zhao et al. (2023b, a). We focus on supervised approaches. Wang et al. (2023) provide the evaluation results based on several supervised detectors, including RoBERTa Liu et al. (2019), XLM-R Conneau et al. (2019), logistic regression classifier with GLTR features Gehrmann et al. (2019b), stylistic features Li et al. (2014), andNELA Horne et al. (2019) features. Similar work using supervised methods can also be seen in Guo et al. (2023); Hu et al. (2023); Xiong et al. (2024). However, most of them only consider binary detection.
@@
The fine-grained multi-class classification problem is closely related to authorship attribution. Munir et al. (2021) find that texts generated from LMs contain distinguishable signals that can be used to attribute the source of texts. Uchendu et al. (2020) investigate three authorship attribution problems: (1) whether two texts are generated by the same method, (2) whether a text is generated by machine or human, and (3) which language model generated the texts. Venkatraman et al. (2023) examine if the principle that humans prefer to spread information evenly during language production can help capture the unique signature of each LLM and human author while Rivera Soto et al. (2024) leverages representations of writing styles.
@@
goes beyond the binary/multi-class classification lens and considers an adversarial setting of co-authorship of human and machine Macko et al. (2024). It is known that machine-generated detection methods are susceptible to authorship obfuscation attacks such as paraphrasing, back-translation, and human-machine collaboration mixture Crothers et al. (2022); Krishna et al. (2023); Shi et al. (2023); Koike et al. (2023). Gao et al. (2024) introduces a dataset with a mixture of machine and human written texts using operations such as polishing, completing Xie et al. (2023), rewriting Shu et al. (2023), adding natural noise Wang et al. (2021), and adapting Gero et al. (2022). Kumarage et al. (2023) use stylometric signals to quantify changes in tweets and detect if and when AI starts to generate tweets. In this work, we focus on detecting and localizing the change point from human-written segments to machine-generated segments. We construct a dataset consisting of peer reviews and student essays based on ChatGPT, GPT-4, LLaMA-2 (7B, 13B and 70B). Models using semantic features show promising results when training and test sets are from the same domain and generators.
@@ 3 Datasets and Metrics
| Domain | License | Human | Human | davinci-003 | ChatGPT | Jais-30B | LLaMA-2-70B | Total | |
@@
The task 1 aims to distinguish human-written from machine-generated text — a binary classification task. Based on the M4 dataset (Wang et al., 2023), given a domain, we up-sampled human text to match the total number of machine-generated text to avoid data imbalance. After upsampling, the dataset consists of 65,177 human-written texts and 73,288 machine-generated texts as shown in Table 1. We additionally generated texts using GPT-4 for each domain to evaluate the detectors’ generalization ability on unseen and strong generators. For multilingual track, we integrate new languages (German and Italian), and news for Arabic by generators of ChatGPT and Jais-30B, as the statistics shown in Table 2. Data Cleaning: simple artifacts in human-written texts, such as multiple newlines and bullet points, could assist detectors, as their presence in the training data may discourage detectors from learning more generalized signals. Therefore, we performed minimal cleaning of the upsampled human-written texts: (i) in a WikiHow text, we removed multiple commas at the beginning of a new line (like ‘‘,,,,,,,,,,, we believe that …’’) and repeating newlines (‘‘nnnnn text begin nnnnn”); (ii) in Wikipedia articles, we removed references (e.g., [1], [2]), URLs, multiple newlines, as well as paragraphs whose length was less than 50 characters; and (iii) in arXiv abstracts, we removed newlines stemming from PDF conversion. Metrics: in addition to accuracy, we also use precision, recall, and F1-score with respect to MGT.
@@ 3.2 Multi-way Detection
Task 2 is to determine which generator model produces the considered text. Different from task 1, we use the parallel subset for task 2 without the upsampling of human-written texts, including six generators: ChatGPT, davinci-003, GPT-4, Cohere, Dolly-v2, and BLOOMz. A new domain OUTFOX is collected to evaluate the classifier’s domain generalization in student essays. Metrics: F1-score, precision, recall in terms of seven individual labels representing different generators and human-written texts are used. We also provide F1-macro and accuracy.
@@ 3.3 Boundary Identification
We consider two common misuse scenarios of human-machine mixed text — continuing to write academic paper reviews and student essays. Given the partial human-written essay with the corresponding problem statement and the partial human-written paper reviews with the title and abstract, we complete them by generating text based on GPT and LLaMA-2 series using prompts as shown in Figure 2. The proportion of words that are human-written range from 0 to 50%. The goal of a detection model is to identify the boundary position from which the text is generated by machines. Mixed texts are specifically generated using ChatGPT, GPT-4, and LLaMA-2 (7B, 13B and 70B). We generate 5,676 and 1,000 examples for two domains respectively using each generator (full statistics is presented in Table 3). Metrics: the Mean Absolute Error (MAE) is used to evaluate the performance of the boundary detection model. It measures the average absolute difference between the predicted position index and the actual changing point.
@@ 4 Human Evaluation
4.1 Data for Human Evaluation
@@ 5 Experiments
We applied two detectors for boundary identification based on semantic features. Five baseline classifiers suggested by Wang et al. (2023) for tasks 1 and 2 include: fine-tuned RoBERTa and XLM-R classifiers, logistic regression with GLTR features (LR-GLTR), SVM with Stylistic (Stylistic-SVM), and NELA features (NELA-SVM).
RoBERTa and XLM-R
We fine-tune two Transformer-based models for MGT detection: RoBERTa and XLM-R.
@@
are based on the assumption that in order to generate fluent and natural-looking text, most LLM decoding strategies sample high-probability tokens from the head of the distribution Gehrmann et al. (2019a). Thus, word ranking information of an LM can be used to distinguish machine-generated texts from human-written. Following Wang et al. (2023), we select two categories of these features: (1) the number of tokens in the top-10, top-100, top-1000, and 1000+ ranks from the LM predicted probability distributions (4 features); and (2) the Frac(p) distribution over 10 bins ranging from 0.0 to 1.0 (10 features). Frac(p) describes the fraction of probability for the actual word divided by the maximum probability of any word at this position. We train a logistic regression model based on these 14 features to perform the binary classification task with the parameter of the maximum iteration = 1,000.
Stylistic and NELA features.
@@
| Detector | Test | Prec | Recall | F1-score | Acc |
@@
For task 1, we simulate a practical scenario where a detector have to deal with texts generated from a new LLM while it is trained on the outputs from a limited number of other generators. Practical detector should be able to generalize over different generators and be robust on unseen inputs.
Experiment Setup
We conduct experiments in two setups. (1) Given a single LLM, we combine its generations across all domains and the upsampled human-written texts as the test set. Texts produced by other LLMs and humans compile training and development sets (split by 4:1). (2) Over all data, we randomly split train and test data across all generators (row All in Table 5). We train detectors with five different seeds for all experiments and present the mean values of the metrics.
@@
Classification results of five detectors are presented in Table 5. We compare results from three aspects: (1) given a detector, on which unseen generator, it generalizes the best (bold F1-score and Acc); (2) given a setting that training without generations of the generator to test, which detector performs the best (underlined Acc); and (3) on average across different settings, which detector is the most accurate and has the best generalization capabilities? Which detector has higher recall than precision? RoBERTa, XLM-R, and LR based on stylistic features give the best F1-score and accuracy on the setting where the classifier is trained with generations of all generators and tested across all generators as well. GLTR features perform best in the setting in which the detector is trained over data without generations from Cohere, and tested over Human vs. Cohere data, while NELA has the highest accuracy in distinguishing human vs. GPT-4 when it has not seen any data from GPT-4. The results for RoBERTa and XLM-R suggest that such detectors encountered notable challenges when confronted with text from new generators not encompassed in their training dataset, demonstrating the large accuracy gap between the setting of All and others. This problem is especially conspicuous in the case of BLOOMz for all detectors (All vs. BLOOMz), followed by Dolly-v2, davinci-003, ChatGPT, GPT-4 and Cohere. XLM-R and GLTR demonstrate better generalization performance (underlines numbers) over (davinci-003, Dolly-v2, and BLOOMz) and (ChatGPT, GPT-4, and Cohere) respectively. XLM-R is the best on average. It has higher accuracy and better generalization ability. LR with NELA features is the worst. Recall of RoBERTa, XLM-R, and GLTR is higher than precision, while the opposite for stylistic and NELA features. In the scenario of detecting misuse, we expect higher recall to avoid less misses of potentially malice.
5.2 Multilingual Human vs. Machine
Results and Analysis
@@
| RoBERTa | All | 96.35 | 96.21 | 96.23 | 96.25 |
| arXiv | 60.55 | 41.11 | 33.67 | 41.11 | |
| PeerRead | 67.54 | 65.47 | 61.55 | 64.77 | |
| 73.59 | 67.97 | 62.33 | 67.97 | ||
| WikiHow | 70.94 | 67.60 | 64.91 | 64.20 | |
| Wikipedia | 40.44 | 50.10 | 37.61 | 49.64 | |
| XLM-R | All | 87.69 | 86.26 | 86.03 | 86.36 |
| arXiv | 51.40 | 41.73 | 38.82 | 41.73 | |
| PeerRead | 52.20 | 51.66 | 44.84 | 50.38 | |
| 69.82 | 60.58 | 59.92 | 60.58 | ||
| WikiHow | 64.01 | 60.81 | 57.38 | 56.13 | |
| Wikipedia | 58.57 | 41.07 | 37.92 | 40.25 | |
| GLTR-LR | All | 42.9 | 46.11 | 43.07 | 45.84 |
| arXiv | 21.5 | 30.61 | 23.19 | 30.60 | |
| PeerRead | 43.50 | 46.24 | 42.52 | 46.37 | |
| 46.88 | 47.91 | 43.90 | 47.91 | ||
| WikiHow | 43.87 | 38.78 | 38.28 | 37.23 | |
| Wikipedia | 41.91 | 35.66 | 35.54 | 34.08 | |
| GLTR-SVM | All | 53.19 | 53.47 | 52.50 | 53.75 |
| arXiv | 27.88 | 35.40 | 29.25 | 35.40 | |
| PeerRead | 33.98 | 37.18 | 34.27 | 37.54 | |
| 46.68 | 47.84 | 44.73 | 47.84 | ||
| WikiHow | 43.88 | 42.36 | 40.52 | 37.84 | |
| Wikipedia | 43.62 | 37.96 | 35.69 | 37.18 | |
| Stylistic-SVM | All | 78.57 | 37.80 | 47.86 | 35.53 |
| arXiv | 46.57 | 7.67 | 11.55 | 7.44 | |
| PeerRead | 53.40 | 27.16 | 30.39 | 25.30 | |
| 61.96 | 28.23 | 33.07 | 26.53 | ||
| WikiHow | 54.48 | 34.89 | 39.94 | 26.49 | |
| Wikipedia | 51.40 | 23.56 | 29.54 | 17.78 | |
| NELA-SVM | All | 66.99 | 27.48 | 34.80 | 25.60 |
| arXiv | 53.46 | 12.97 | 19.11 | 11.71 | |
| PeerRead | 48.53 | 24.76 | 22.84 | 23.31 | |
| 45.94 | 28.09 | 31.45 | 24.20 | ||
| WikiHow | 46.87 | 24.81 | 28.54 | 21.16 | |
| Wikipedia | 47.81 | 22.62 | 27.36 | 19.79 | |

### v2

@@
M4GT-Bench: Evaluation Benchmark
for Black-Box Machine-Generated Text Detection
@@
M4GT-Bench: Evaluation Benchmark
for Black-Box Machine-Generated Text Detection
Yuxia Wang,† Jonibek Mansurov,† Petar Ivanov,† Jinyan Su,† Artem Shelmanov,† Akim Tsvigun,† Osama Mohammed Afzal,† Tarek Mahmoud,† Giovanni Puccetti,§ Thomas Arnold,¶ Alham Fikri Aji,† Nizar Habash,†‡ Iryna Gurevych,† Preslav Nakov† †Mohamed bin Zayed University of Artificial Intelligence, UAE ¶TU Darmstadt, Germany §Institute of Information Science and Technology, Italy ‡New York University Abu Dhabi, UAE {yuxia.wang, jonibek.mansurov, preslav.nakov}@mbzuai.ac.ae
@@
M4GT-Bench: Evaluation Benchmark
for Black-Box Machine-Generated Text Detection
Yuxia Wang,† Jonibek Mansurov,† Petar Ivanov,† Jinyan Su,† Artem Shelmanov,† Akim Tsvigun,† Osama Mohammed Afzal,† Tarek Mahmoud,† Giovanni Puccetti,§ Thomas Arnold,¶ Alham Fikri Aji,† Nizar Habash,†‡ Iryna Gurevych,† Preslav Nakov† †Mohamed bin Zayed University of Artificial Intelligence, UAE ¶TU Darmstadt, Germany §Institute of Information Science and Technology, Italy ‡New York University Abu Dhabi, UAE {yuxia.wang, jonibek.mansurov, preslav.nakov}@mbzuai.ac.ae
@@ 1 Introduction
Task 2: Multi-Way Machine-Generated Text’s Generator Detection. This task involves identifying the specific generator responsible for producing a given text. The text may be either human-written or crafted by a LLM. A different perspective to look at Task 2 is that we are attributing authorship to a specific generator Munir et al. (2021).
Task 3: Human-Written to Machine-Generated Text Change Point Detection. In this task, the goal is to precisely recognize the boundary within the mixed text, where the transition occurs from human-written to machine-generated content.
@@
We construct a diverse public multilingual MGT detection benchmark involving nine languages, six domains, nine LLM generators and three different tasks.
We introduce a novel task formulation for MGT detection, where human-machine mixed text is explored. It is the first attempt to evaluate the ability of automatic approaches on detecting a boundary between human-written and LLM-generated texts.
On the developed benchmark, we have tested several strong MGT detection baselines and also conducted an evaluation of human performance. We see that obtaining good performance in MGT detection usually requires an access to the training data from the same domain and generators.
@@ 2 Related Work
The task of detecting MGTs has traditionally been formulated as a binary classification problem Zellers et al. (2019); Gehrmann et al. (2019a); Solaiman et al. (2019); Ippolito et al. (2019). MGT detection can be broadly categorized into two main types: supervised and unsupervised methods. Supervised approaches Wang et al. (2023); Uchendu et al. (2021); Zellers et al. (2019); Zhong et al. (2020); Liu et al. (2022) rely on annotated datasets to train classifiers. In contrast, unsupervised methods leverage white-box features such as likelihood and log-rank Solaiman et al. (2019); Ippolito et al. (2019); Mitchell et al. (2023); Su et al. (2023); He et al. (2023); Hans et al. (2024), or employ watermarking techniques Kirchenbauer et al. (2023); Zhao et al. (2023b, a) to identify machine-generated text. Here, we focus on supervised approaches. Wang et al. (2023) evaluate several supervised detectors, including RoBERTa Liu et al. (2019), XLM-R Conneau et al. (2019), logistic regression classifier with GLTR features Gehrmann et al. (2019b), stylistic features Li et al. (2014), and NELA Horne et al. (2019) features. Similar investigation of supervised methods is also done by Guo et al. (2023); Hu et al. (2023); Xiong et al. (2024).
@@
There is a growing interest for more fine-grained classification that can not only identify the nature of the texts (i.e., whether it is machine-generated or human written), but also its specific source (i.e., which LLM generates it?). The fine-grained multi-class classification problem is closely related to the authorship attribution Uchendu et al. (2020); Venkatraman et al. (2023); Rivera Soto et al. (2024). Munir et al. (2021) find that texts generated by LLMs contain distinguishable signals that can be used to attribute the source of texts. Uchendu et al. (2020) investigate three authorship attribution problems: (1) whether two texts are produced by the same generator, (2) whether a text is generated by a machine or a human, and (3) which LLM generated the text. Venkatraman et al. (2023) examine if the principle that humans prefer to spread information evenly can help capturing unique signatures of LLMs and human authors. Similarly, Rivera Soto et al. (2024) leverage representations of writing styles.
@@
Work on authorship obfuscation extends beyond the binary/multi-class classification perspective by addressing the adversarial context of co-authorship between human and machine Macko et al. (2024). Several works show that MGT detection methods are susceptible to authorship obfuscation attacks such as paraphrasing, back-translation, and human-machine collaboration mixture Crothers et al. (2022); Krishna et al. (2023); Shi et al. (2023); Koike et al. (2023). Gao et al. (2024) introduces a dataset with a mixture of machine and human written texts using operations such as polishing, completing Xie et al. (2023), rewriting Shu et al. (2023), adding natural noise Wang et al. (2021), and adapting Gero et al. (2022). Kumarage et al. (2023) use stylometric signals to quantify changes in tweets and detect if and when AI starts to generate tweets. Dugan et al. (2020, 2023) investigate human ability to detect the boundary, where machine generated continuations are from LLMs. However, they use currently outdated LLMs such as GPT2-XL. In this work, we suggest the task of automatic detection and localization of change points from human-written segments to machine-generated segments. Our task is more challenging than previous works as we used state-of-the-art LLMs for generating continuations.
@@ 3 Datasets and Metrics
Domain | License | Human | Human | davinci-003 | ChatGPT | Jais | LLaMA-2 | Total | |
@@
The task 1 aims to distinguish human-written from machine-generated text — a binary classification task. Based on the M4 dataset (Wang et al., 2023), given a domain, we up-sampled human text to match the total number of machine-generated text to avoid data imbalance. After upsampling, the dataset consists of 65,177 human-written texts and 73,288 machine-generated texts as shown in Table 1. We additionally generated texts using GPT-4 for each domain to evaluate the detectors’ generalization ability on unseen and strong generators. For the multilingual track, we integrate new languages (German and Italian), and news for Arabic by generators of ChatGPT and Jais-30B, as the statistics shown in Table 2.
Data Cleaning
Simple artifacts in human-written texts, such as multiple newlines and bullet points, could assist detectors. Their presence in the training data may discourage detectors from learning more generalized signals. Most importantly, these artifacts primarily originate from data crawling (multiple newlines at the beginning and the end of text) and conversion (such as PDF to text). To recover the original format of human-written plain text, we performed minimal cleaning on the upsampled human-written texts: (i) in a WikiHow text, we removed multiple commas at the beginning of a new line (e.g. ‘‘,,,,,,, we believe that …’’) and repeating newlines (‘‘nnnn text begin nnnn”); (ii) in Wikipedia articles, we removed references (e.g., [1], [2]), URLs, multiple newlines, as well as paragraphs whose length was less than 50 characters; and (iii) in arXiv abstracts, we removed newlines stemming from PDF conversion.
Metrics
In addition to accuracy, we also use precision, recall, and F1-score with respect to MGT.
@@ 3.2 Multi-way Detection
Task 2 is to determine which generator model produces the considered text. This is motivated by the Intelligence Protection (IP) of LLMs generations. For example, it can be used to detect whether model A generations are extensively used to train another commercial model B by distillation (this is forbidden by OpenAI). If many generations from B are detected as model A generations, one might suspect that model B used a large amount of generations of model A in the training process. In general, identifying the specific model used to produce a copyrighted material is important for legal and copyright reasoning. Additionally, in cases where the generated material is harmful, misleading, or illegal, pinpointing the exact LLM responsible is essential for addressing ethical concerns and legal obligations. Different from task 1, we use the parallel subset for task 2 without the upsampling of human-written texts, including six generators: ChatGPT, davinci-003, GPT-4, Cohere, Dolly-v2, and BLOOMz. A new domain OUTFOX is collected to evaluate the classifier’s domain generalization in student essays.
Metrics
F1-score, precision, recall in terms of seven individual labels representing different generators and human-written texts are used. We also provide F1-macro and accuracy.
@@ 3.3 Boundary Identification
Task 3 aims to imitate the real-world LLM usage situation, where not the full text is generated by machine, but partially written by human and partially generated by a machine. To simplify the task, we formulate it as first written by a human and then continued by a machine, and the task is to detect the single boundary of change. We consider two common misuse scenarios of human-machine mixed text — continuing to write academic paper reviews and student essays. Given the partial human-written essay with the corresponding problem statement and the partial human-written paper reviews with the title and abstract, we complete them by generating text based on GPT and LLaMA-2 series using prompts as shown in Figure 2. The proportion of words that are human-written range from 0 to 50%. The goal of a detection model is to identify the boundary position from which the text is generated by machines. Mixed texts are specifically generated using ChatGPT, GPT-4, and LLaMA-2 (7B, 13B and 70B). We generate 5,676 and 1,000 examples for two domains respectively using each generator (full statistics is presented in Table 3).
How realistic is the task setting?
Providing the beginning of a text and asking an LLM to generate the rest is a common use case. At the same time, we believe that MGTs seriously modified by humans might not represent a malicious use case of LLMs, as LLMs are legitimate writing tools. Depending on the amount of human intervention, one could argue that this is not anymore machine-generated, but machine-human collaboration. Therefore, we seek balance between the complexity of the task and its practicality. Moreover, detecting texts that contain multiple changing points is an interesting research direction for future work, but it would bring new challenges with analyzing shorter spans. Extremely, if a human added an extra word in a machine-generated sentence, it would be much harder for blackbox detectors to identify this word.
Metrics
Mean Absolute Error (MAE) is used to evaluate the boundary detection model’s performance. It measures the average absolute difference between the predicted position index and the actual changing point.
@@ 4 Human Evaluation
Data for Human Evaluation
@@ 5 Experiments
We conduct an experimental evaluation with several common baselines representing a diverse set of detection techniques: standard fine-tuning of an encoder-based Transformer model, elaborated feature engineering for the similar task, ranking-based features which form patterns for different LLMs. For classification tasks 1 and 2, we use five baseline classifiers suggested by Wang et al. (2023). They include fine-tuned RoBERTa and XLM-R classifiers, logistic regression with GLTR features (LR-GLTR) Gehrmann et al. (2019a), SVM with Stylistic (Stylistic-SVM), and SVM with NELA features (NELA-SVM) (Horne et al., 2019). For the boundary detection task 3, we applied two detectors based on semantic features.
@@
These features are based on the assumption that in order to generate fluent and natural-looking text, most LLM decoding strategies sample high-probability tokens from the head of the distribution Gehrmann et al. (2019a). Thus, word ranking information of an LM can be used to distinguish machine-generated texts from human-written. Following Wang et al. (2023), we select two categories of these features: (1) the number of tokens in the top-10, top-100, top-1000, and 1000+ ranks from the LM predicted probability distributions (4 features); and (2) the Frac(p) distribution over 10 bins ranging from 0.0 to 1.0 (10 features). Frac(p) describes the fraction of probability for the actual word divided by the maximum probability of any word at this position. We train a logistic regression model based on these 14 features to perform the binary classification task with the parameter of the maximum iteration = 1,000.
Stylistic and NELA features
@@
Detector | Test | Prec | Recall | F1 | Acc |
@@
For Task 1, we simulate a practical scenario where a detector have to deal with texts generated from a new LLM while it is trained on the outputs from a limited number of other generators. A practical detector should be able to generalize over different generators and be robust on unseen inputs.
Experimental Setup
We conduct experiments in two setups. (1) In the first setup, we select an LLM and combine its generations across all domains with the upsampled human-written texts to create a test set. Texts produced by other LLMs and humans are divided into training and development sets with a 4:1 ratio. (2) In the second setup, we take the data from all domains and generators and randomly split it into train and test sets (row All in Table 5). We train detectors with five different seeds for all experiments and present the mean values of the metrics.
@@
Classification results of five detectors are presented in Table 5. We analyze the results to answer three research questions: (1) For a given detector, on which unseen generator does it generalize best (highlighting F1-score and accuracy)? (2) Given a setting that training without the data of the generator to test, which detector performs the best (underlined Acc). (3) On average across both settings, which detector is the most accurate and has the best generalization capabilities? Which detector has higher recall than precision? RoBERTa, XLM-R, and LR based on stylistic features give the best F1-score and accuracy on the setting where the classifier is trained with generations of all generators and tested across all generators. GLTR features perform the best in the setting where the detector is trained on data without generations from Cohere, and tested over Human vs. Cohere data, while NELA has the highest accuracy in distinguishing human vs. GPT-4 when it has not seen any data from GPT-4. The results for RoBERTa and XLM-R suggest that such detectors encountered notable challenges when confronted with text from new generators not encompassed in their training data, exhibiting the large accuracy gap between the setting of All and others. This problem is especially conspicuous in the case of BLOOMz for all detectors (All vs. BLOOMz), followed by Dolly-v2, davinci-003, ChatGPT, GPT-4, and Cohere. XLM-R and GLTR demonstrate better generalization performance (underlines numbers) over (davinci-003, Dolly-v2, and BLOOMz) and (ChatGPT, GPT-4, and Cohere) respectively. XLM-R is the best on average. It has higher accuracy and better generalization ability. LR with NELA features is the worst. Recall of RoBERTa, XLM-R, and GLTR is higher than precision, while it is the opposite for stylistic and NELA features. In the scenario of detecting misuse, we expect higher recall to avoid less misses of potentially malice. Overall, detectors show poor generalization performance when testing on unseen generators. This is well-known due to exposure bias. The generations by BLOOMz are significantly different from other generators, leading to the result that when training on the data combination of other generators and testing on BLOOMz, the performance is extremely low for all detectors. The detector based on NELA features overall performs the worst, which demonstrates that the features used for fact verification are not suitable to serve as signals to distinguish human vs. machine text.
@@
RoBERTa | All | 96.96 | 97.01 | 96.94 | 97.00 |
arXiv | 55.72 | 36.55 | 32.29 | 36.55 | |
PeerRead | 70.58 | 70.12 | 66.89 | 69.47 | |
77.21 | 74.49 | 71.66 | 74.49 | ||
WikiHow | 72.62 | 70.56 | 68.85 | 68.36 | |
Wikipedia | 46.37 | 52.22 | 39.37 | 51.91 | |
OUTFOX | 71.71 | 65.04 | 66.40 | 78.25 | |
XLM-R | All | 90.73 | 90.37 | 90.16 | 90.17 |
arXiv | 51.29 | 43.88 | 41.71 | 43.88 | |
PeerRead | 53.68 | 52.14 | 46.10 | 50.91 | |
69.73 | 58.72 | 57.34 | 58.71 | ||
WikiHow | 65.73 | 60.84 | 58.45 | 57.38 | |
Wikipedia | 60.04 | 42.55 | 38.80 | 41.95 | |
OUTFOX | 51.94 | 42.44 | 43.00 | 52.10 | |
GLTR-LR | All | 42.36 | 43.96 | 40.32 | 45.06 |
arXiv | 26.24 | 34.45 | 26.92 | 34.45 | |
PeerRead | 42.20 | 44.10 | 39.04 | 44.32 | |
45.54 | 46.28 | 41.7 | 46.28 | ||
WikiHow | 41.86 | 39.39 | 38.24 | 38.74 | |
Wikipedia | 41.62 | 36.95 | 34.62 | 35.18 | |
OUTFOX | 29.68 | 32.38 | 29.18 | 38.78 | |
GLTR-SVM | All | 52.81 | 48.42 | 47.24 | 50.39 |
arXiv | 22.57 | 32.29 | 25.73 | 32.28 | |
PeerRead | 34.10 | 39.10 | 34.81 | 39.40 | |
43.21 | 46.19 | 40.94 | 46.19 | ||
WikiHow | 44.13 | 39.01 | 38.27 | 36.18 | |
Wikipedia | 34.86 | 30.95 | 26.72 | 29.81 | |
OUTFOX | 26.93 | 28.29 | 26.45 | 28.58 | |
Stylistic-SVM | All | 78.95 | 37.10 | 47.31 | 35.26 |
arXiv | 44.89 | 8.52 | 12.71 | 8.27 | |
PeerRead | 50.72 | 21.96 | 25.43 | 20.44 | |
60.98 | 27.25 | 31.16 | 24.43 | ||
WikiHow | 56.23 | 34.57 | 37.04 | 26.71 | |
Wikipedia | 47.94 | 21.21 | 27.77 | 16.06 | |
OUTFOX | 48.28 | 27.46 | 32.50 | 26.80 | |
NELA-SVM | All | 64.50 | 23.91 | 30.54 | 22.76 |
arXiv | 47.35 | 11.53 | 16.20 | 10.94 | |
PeerRead | 44.63 | 20.00 | 20.97 | 18.72 | |
42.77 | 24.27 | 27.87 | 20.72 | ||
WikiHow | 48.33 | 25.81 | 25.51 | 21.83 | |
Wikipedia | 46.38 | 20.74 | 25.06 | 18.76 | |
OUTFOX | 35.05 | 17.44 | 18.48 | 19.18 |